OpenStudy (anonymous):

PLEASE HELP The family trip to Grandma's consisted of both a train ride and a car ride. The average speed of the train ride was 72 miles per hour, and the average speed of the car ride was 51 miles per hour. The entire trip lasted 10 hours. Let x be the number of hours the train ride lasted. Write an expression for the total distance of the trip, in miles.

OpenStudy (gorv):

distance travelled by train = speed *time =72*x=72x time by car=10-x distance by car=51*(10-x)=510-51x total distance=72x+510-51x =21x+510